Evaluate the given expression and express the result using the usual format for writing numbers (instead of scientific notation)
57P2
this is what i tried 57/(57-2)= 57/55

Answer by stanbon(75887)

57P2
-----
= 57!/(57-2)! = 57!/55! = 57*56 = 3192
----------
Cheers,
Stan H.